OpenStudy (lgbasallote):

An OpenStudy user's speed is 40 medals per day and his acceleration is 2 medal/day^2. If it takes 4,500 medals to reach the Legend rank, how many days will it take for him to reach the rank?

OpenStudy (anonymous):

lol in first

OpenStudy (anonymous):

wrong group

OpenStudy (lgbasallote):

isnt it math?

OpenStudy (anonymous):

What is his/her initial medalage? does he/she start from 0?

OpenStudy (anonymous):

physics

OpenStudy (lgbasallote):

yup 0

OpenStudy (lgbasallote):

but isnt this an ordinary math problem too?

OpenStudy (lgbasallote):

i mean the basic concepts still come in play

OpenStudy (anonymous):

its both, physics and math.Arguably physics is nothing without math though :)

OpenStudy (lgbasallote):

lol can we just teach me how to solve this? :p

OpenStudy (anonymous):

well first of all, his acceleration cant be 2 medals/day a = v / t since you postulated that v = 40 per day a = 40 per day because a = delta V / delta t

OpenStudy (lgbasallote):

but my book says 50 days o.O

OpenStudy (anonymous):

I would consider it an initial velocity. Its like saying a car is going at 40 ft/s, and accelerating at 2 ft/s^2.

OpenStudy (lgbasallote):

yeah that'swhat i think too

OpenStudy (anonymous):

To solve the problem use the formula:\[d = x_0+v_0t+\frac{1}{2}at^2\]Set d to 4500, intial position to 0, intiial velocity to 40, and acceleration to 2. Solve for t.

OpenStudy (lgbasallote):

so 4500 = 40t + t^2?

OpenStudy (anonymous):

yep

OpenStudy (lgbasallote):

hmm yep i think that gives 50thanks